### v1

# An intense and controlled beam of S() atoms

Abstract
We report the production of an intense and controlled beam of electronically excited sulfur atoms in the state using a multistage Zeeman decelerator. Sulfur atoms, generated via photolysis of CS2, are produced in both the ground and excited states. We demonstrate that both can be manipulated using the decelerator, and that temporal separation between them can be achieved by operating in deceleration mode. This enables the generation of sulfur atom beams with a well-defined velocity, narrow velocity spreads, and an enhanced quantum-state purity. To assess the suitability of the beam for scattering studies, we performed a proof-of-principle elastic collision experiment with S() and argon atoms. The observed velocity-map-imaging signal confirms that the S() beam density is sufficient for detailed scattering studies. These results form the foundation for future studies of reactive and quenching processes involving S() atoms at tunable and well-defined collision energies.
I Introduction
Sulfur (S) atoms play an important role in a wide range of chemical environments, from combustion and atmospheric processes to interstellar chemistry. In their electronically excited state, S atoms can undergo electronic quenching, such that they relax to the ground state via collisions [1, 2, 3, 4, 5], or participate in reactive collisions that lead to the formation of new chemical products [1, 6, 7, 8, 9, 10]. These competing pathways offer valuable insights into fundamental energy transfer mechanisms and chemical reactivity. In particular, the S() + H2 reaction serves as one of the prototypical systems for investigating the dynamics of complex-forming or insertion reactions [10, 11, 12, 13, 14].
To investigate such processes experimentally, the crossed molecular beam technique has proven to be a powerful tool, providing detailed insights into the underlying reaction dynamics and energy transfer mechanisms [15, 16, 17, 18, 19, 20, 21, 22, 23]. In these experiments, velocity map imaging (VMI) is commonly used to detect the angular and speed distributions of the collision products in order to extract differential cross sections [24]. Generally, the resolution in these experiments, both in terms of collision energy and image resolution, is limited by the properties of the atomic and molecular beams involved. Broad velocity distributions as well as insufficient quantum-state control can obscure subtle features in the scattering dynamics and complicate comparisons with theoretical predictions.
To overcome these limitations and to study sulfur atom collisions in the highest possible detail, it would be favorable to produce an intense and controlled beam of S() atoms with a well-defined velocity, narrow velocity spreads, and, when necessary, a high quantum-state purity. Such a beam would allow for precise control over the collision energy and could significantly improve the resolution of VMI images, thereby enhancing the level of detail accessible in scattering experiments.
In previous crossed-beam scattering studies, beams of S() atoms have been generated by ultraviolet (UV) photolysis of supersonically expanded carbon disulfide (CS2), typically seeded in helium or neon [25, 26, 27, 28, 29, 30, 31]. This photolysis process produces S atoms in both the excited and ground states, with branching ratios that strongly depend on the photolysis wavelength [32, 33, 34, 35, 36, 37, 38, 39, 40]. Most scattering studies have employed 193 nm radiation, which is readily available from an ArF excimer laser, [25, 26, 27, 31] while others used tunable radiation near 195 nm [28, 29, 30] to produce S() .
In this work, we demonstrate the production of a high-intensity, velocity-controlled beam of electronically excited S() atoms using a 3-m-long multistage Zeeman decelerator. The sulfur atoms are generated via photolysis of CS2 at 199.65 nm, which efficiently produces both S() and ground-state S() atoms. We show that both can be manipulated using the decelerator, and that temporal separation between these states can be achieved by decelerating the S atoms. The resulting S() beam has a well-defined velocity, narrow velocity spreads, and a high quantum-state purity if needed, making it ideally suited for high-resolution scattering experiments. As a proof of concept, we present an elastic scattering experiment involving collisions between S() and argon (Ar) atoms. The observed scattering signals confirm that the beam density is sufficient for detailed collision experiments.
II Experiment
The experiments were conducted in a molecular beam apparatus containing a 3-m-long Zeeman decelerator, illustrated in Fig. 1. This machine has been used to decelerate beams of O atoms [41], O2 molecules [41], NO (X ) radicals [42], NH radicals [43], and C atoms [44]. A detailed description of the decelerator can be found elsewhere [45]. Here, we focus on the experimental aspects related to the production, control, and detection of S atoms.
A mixture of 0.5% CS2 seeded in neon with a backing pressure of 3 bar was expanded through a Nijmegen pulsed valve (NPV) [46], operated at a repetition rate of 20 Hz. At the exit of the valve, S atoms were produced by photolysis of CS2 using a pulse energy of 0.45 mJ in a 5 ns pulse of 199.65 nm tunable radiation, focused using a cylindrical lens with 400 mm focal length. This UV radiation was obtained by frequency tripling the output of a dye laser (LIOP-TEC), pumped using the second harmonic of an Nd:YAG laser (InnoLas SpitLight 600). High S-atom signals were found at this wavelength, in agreement with Ref. [33]. The S atoms with a mean forward velocity of around 875 m/s were formed in both the ground as well as the excited states. In the ground state, the sulfur atoms have a magnetic moment of 3, 1.5, and 0 for the spin-orbit levels 2, 1, and 0, respectively. The S atoms in the excited state have a magnetic moment of 2 . In the presence of a magnetic field, these states split into the different components, where is the projection of the total electronic angular momentum on the space-fixed -axis. The resulting Zeeman energy-level diagrams are shown in Fig. 2. Sulfur atoms that reside in low-field-seeking components, depicted in red in Fig. 2, can be selected by the Zeeman decelerator. Atoms in the components are insensitive to magnetic fields and pass through the decelerator in free flight, which strongly reduces their density and thus their contribution.
Approximately 100 mm downstream from the nozzle of the NPV, the S atoms passed a 3-mm-diameter skimmer and entered the multistage Zeeman decelerator [45], consisting of an alternating array of 127 pulsed solenoids and 127 permanent hexapoles. Afterwards, the beam was guided towards the detection region using 7 additional hexapoles.
To record time-of-flight (TOF) profiles, the S atoms were detected state-selectively using resonance-enhanced multiphoton ionization (REMPI). The S() products were ionized by (1+1) REMPI via the (4s) state at 216.9 nm [47] using a pulse energy of 0.45 mJ in a 5 ns pulse. The S() products were ionized by (2+1) REMPI via the (4p) states using wavelengths between 308 and 311 nm [48], using a pulse energy of 0.14 mJ in a 5 ns pulse. The first laser intersected the molecular beam at , and the second at . In both cases, the laser beam was focused in the interaction region by a spherical lens with a focal length of 500 mm. Both laser beams were generated by Nd:YAG (InnoLas SpitLight 1500) pumped dye lasers (LIOP-TEC).
The resulting S ions were accelerated by high-resolution VMI ion optics [49] towards a mass-gated microchannel plate detector. The impact positions of the ions were recorded using a phosphor screen in combination with a CMOS camera.
III Results and discussion
III.1 TOF measurements
Within our experimental framework, the decelerator was used in two modes of operation, as described in Ref. [45]. In deceleration mode, each solenoid was pulsed once, leading to a nearly constant reduction of kinetic energy per solenoid and a corresponding decrease in the longitudinal velocity of the S-atom packet. In the so-called hybrid mode, each solenoid was pulsed twice to guide the S-atom packet through the decelerator at a constant speed. In both modes, the final velocity of the packet was determined by the specific pulse sequence used for the solenoids.
In this study, we focused on final longitudinal velocities around 875 m/s, which can be achieved by seeding in neon. While effective guiding of S atoms is possible in this case, the relatively low magnetic moment-to-mass ratio of S, in combination with the relatively high velocity, limits the amount of deceleration achievable. Other beam velocities could for instance be reached by using a different seed gas. However, using pure helium results in beam velocities that are too high to be efficiently manipulated by the decelerator. On the other hand, heavier rare gases such as argon or krypton produce slower beams, but are known to have significantly higher quenching cross sections for S() atoms [4], thereby reducing the S() density. Yet, if the beam density is not the main limitation, these heavier seed gasses could be employed to reach lower final velocities.
Representative TOF profiles of S() atoms at various velocities, obtained using the hybrid mode, are shown in blue in the upper part of Fig. 3. Only the segments of the TOF profiles containing the selected packets are displayed. The final velocity of the packet of S() atoms was controlled by selecting a different initial velocity of the beam pulse and guiding this packet at constant speed through the decelerator. For comparison, the TOF profile recorded without pulsing the solenoids, i.e., when the beam is only transversally focussed by the hexapoles, is shown in black. The hybrid mode of the decelerator clearly produces narrow and intense peaks, indicating an effective longitudinal control of the S() atoms.
The measured TOF profiles exhibit good agreement with simulated profiles, shown below the experimental data in Fig. 3. These simulated TOFs result from three-dimensional numerical particle trajectory simulations that account for the forces exerted on the S atoms by the space- and time-dependent magnetic fields inside the Zeeman decelerator [50, 45]. The simulations reproduce both the arrival time distributions of the selected packets and the relative intensities of the observed peaks well. This indicates that the properties of the S-atom beam and its transportation through our machine are well understood.
Since the REMPI scheme used for detecting S() atoms imposes around 12 m/s recoil to the S ions, direct extraction of the velocity distributions of the selected S() atoms from the experimental data is not feasible. Instead, these can be determined from the numerical particle trajectory simulations. For the profiles shown here, the longitudinal velocity spread is around 7.5 m/s, whereas the transverse spread is around 5 m/s, both given as full width at half maximum (FWHM) values.
III.2 Quantum-state distributions
Photolysis of CS2 at 199.65 nm produces S atoms not only in the excited state, but also in the ground states. This is illustrated in the bottom half of Fig. 4, which shows measured REMPI spectra for S() atoms (a) and ground-state S atoms (b) exiting the decelerator with the solenoids switched off. Note that the intensities for ground- and excited-state S atoms are not directly comparable, since different lasers and REMPI schemes were employed. Each nominal transition probing the S() levels contains two or three spectral lines, corresponding to resonant transitions to the different spin-orbit levels of the electronically excited state () [48]. These transitions are indicated by vertical black lines in the figure. Around 87% of the observed ground-state signal originates from the state.
When the decelerator is operated in hybrid mode with a pulse sequence optimized for S() atoms, the signal increases significantly, see the top half of Fig. 4a. Since S() atoms can also occupy low-field-seeking states, see Fig. 2, they are amenable to magnetic manipulation as well, as shown in the top half of Fig. 4b. Comparing the top and bottom halves of Fig. 4b reveals enhanced signals for the and states, even though the pulse sequence was optimized for S() atoms with a magnetic moment of . This increase reflects their interaction with the magnetic fields. In contrast, the intensity remains unchanged, reflecting that this state does not interact with magnetic fields, see Fig. 2. Under these conditions, approximately 91% of the observed ground-state signal can be attributed to the state.
In certain scattering experiments, it is undesirable that the S() and S() atoms arrive simultaneously in the interaction region. For instance, in studies of S() quenching to the ground state, this overlap would lead to contributions of unscattered S() atoms that could obscure part of the scattering signal of interest. The use of a Zeeman decelerator enables temporal separation of these states. Here, we focus on separating the S() from the S() state, which is the most populated ground-state level (see Fig. 4) and possesses the highest magnetic moment (see Fig. 2).
To demonstrate this, we operated the decelerator in two modes to reach the same final velocity of 850 m/s: hybrid mode to guide at a constant velocity, and deceleration mode starting from an initial velocity of 875 m/s. Although both pulse sequences are optimized for selecting S() atoms in the state, other low-field-seeking states of both the and states are transmitted through the decelerator as well. Figs. 5a and b show selected parts of the corresponding TOF profiles recorded for S() (blue) and S() (red) atoms for the hybrid and deceleration mode, respectively. Each profile is normalized to its peak intensity to facilitate comparison.
In hybrid mode, the S() and S() packets exit the decelerator simultaneously, and no separation of atoms in both quantum states is possible. However, in deceleration mode, the S() packet arrives later than the S() packet due to its larger magnetic moment and therefore stronger interaction with the magnetic fields. By selecting the arrival time corresponding to the peak of the S() packet, the quantum-state purity of the packet can be significantly improved. Note that, in a similar way, a high-purity S() packet could be produced as well.
The measured TOF profiles show good agreement with the corresponding simulated profiles, shown in Figs. 5c and d for the S() state and in Figs. 5e and f for S() , although some experimental features are slightly broader or shifted. In all profiles, multiple peaks are observed, originating from different low-field-seeking states with distinct magnetic moments. This is illustrated by the simulated -state-resolved TOF profiles, shown by the dashed and dash-dotted lines in Fig. 5 for and , respectively . The state is not shown, since it has no interaction with the magnetic field and therefore is negligible in intensity. High-field-seeking states are defocused by the hexapoles in the decelerator, and therefore do not reach the detection region.
The level of the state has the highest magnetic moment (). It is therefore decelerated the most and arrives latest at a mean TOF of 3.751 ms (Fig. 5f). The level of the state, with a magnetic moment of , arrives slightly earlier at a mean TOF of 3.737 ms (Fig. 5d). The levels of the and states, with magnetic moments of and , respectively, result in lower-intensity peaks at earlier arrival times due to their lower interaction with the magnetic field. This demonstrates the decelerator’s ability to temporally resolve quantum states of S atoms based on their magnetic moments. It should be noted, however, that in our experimental setup, the sublevels of a given state become degenerate and redistribute before detection due to an insufficient magnetic field between the end of the decelerator and the detection region. As a result, they cannot be distinguished experimentally.
III.3 Elastic scattering
Estimating particle densities from the observed signal levels in our experiment is particularly challenging, and we therefore refrain from this. Instead, to assess whether the beam intensity is sufficient for scattering studies, we performed a proof-of-principle experiment involving elastic collisions between S() and Ar atoms.
In this experiment, the S() atoms were guided through the decelerator in hybrid mode to a final velocity of 875 m/s and intersected a pure Ar beam with a velocity of around 610 m/s at an intersection angle of 45∘. This beam was produced by expanding 3 bar Ar through an Even-Lavie valve [51], operated at a repetition rate of 20 Hz. This resulted in a collision energy of about 285 . Fig. 6 shows the resulting VMI image of the scattered S() atoms, recorded using the 1+1 REMPI detection scheme described in the Experimental section. In the image, the relative velocity vector is oriented horizontally, with forward scattering appearing on the right-hand side. The observed signal levels were sufficient to clearly resolve the scattering distribution with a favorable signal-to-noise ratio. On average, we detected around 1-2 events per laser shot by reducing the laser power to only 55 J, demonstrating that the S() beam produced by our decelerator is intense enough for scattering studies.
IV Conclusions
We have demonstrated the successful production of a high-intensity and velocity-controlled beam of electronically excited S() atoms using our multistage Zeeman decelerator. Following photolysis of CS2, both S() and ground-state S() atoms are formed, and both can be manipulated using the decelerator. By operating the decelerator in deceleration mode, we achieved temporal separation between S() and S() atoms, enabling the generation of a S-atom beam with high quantum-state purity when required.
Furthermore, we showed that the density of the S() beam is sufficient to perform elastic scattering experiments, as evidenced by our VMI image obtained for S() + Ar collisions. This is an excellent starting point for investigating reactive scattering between S() and H2 (including its isotopologues), as well as electronic quenching of S() in collisions with various partners, although it should be noted that the cross sections for these processes are generally expected to be smaller than those for the elastic S() + Ar collisions investigated here. Our experimental setup allows for precise control over the collision energy, making it well suited for exploring these processes under well-defined conditions.
V Acknowledgements
We are grateful to S.Y.T. van de Meerakker for his continued support and for many valuable discussions. We thank X.-D. Wang, S. Kusters, and B. Vonk for their contributions to preparatory studies, and A. van Roij, E. Sweers, and G. Wulterkens for their expert technical support. This work is part of the research program of the Dutch Research Council (NWO). J.O. acknowledges support from the European Research Council (IQ-SCORES, Grant Agreement No. 101160597).

### v2

@@
-An intense and controlled beam of S() atoms
+Intense and controlled beam of S() atoms
 Abstract
@@
 We show that both can be manipulated using the decelerator, and that temporal separation between these states can be achieved by decelerating the S atoms.
-The resulting S() beam has a well-defined velocity, narrow velocity spreads, and a high quantum-state purity if needed, making it ideally suited for high-resolution scattering experiments.
+The resulting S() beam has a well-defined velocity, narrow velocity spreads, and an enhanced quantum-state purity if needed, making it ideally suited for high-resolution scattering experiments.
 As a proof of concept, we present an elastic scattering experiment involving collisions between S() and argon (Ar) atoms.
@@
 Each profile is normalized to its peak intensity to facilitate comparison.
-In hybrid mode, the S() and S() packets exit the decelerator simultaneously, and no separation of atoms in both quantum states is possible.
+In hybrid mode, the S() and S() packets exit the decelerator simultaneously, and it is not possible to separate the atoms in the two quantum states.
 However, in deceleration mode, the S() packet arrives later than the S() packet due to its larger magnetic moment and therefore stronger interaction with the magnetic fields.
 By selecting the arrival time corresponding to the peak of the S() packet, the quantum-state purity of the packet can be significantly improved.
+The purity could be further enhanced by employing a heavier seed gas, which lowers the beam velocity and increases the effectiveness of deceleration, thereby improving the temporal separation between the different states.
+This approach, however, would also reduce the overall S() density, as discussed in section III.1.
 Note that, in a similar way, a high-purity S() packet could be produced as well.
@@
 Following photolysis of CS2, both S() and ground-state S() atoms are formed, and both can be manipulated using the decelerator.
-By operating the decelerator in deceleration mode, we achieved temporal separation between S() and S() atoms, enabling the generation of a S-atom beam with high quantum-state purity when required.
+By operating the decelerator in deceleration mode, we achieved temporal separation between S() and S() atoms, enabling the generation of a S-atom beam with enhanced quantum-state purity when required.
 Furthermore, we showed that the density of the S() beam is sufficient to perform elastic scattering experiments, as evidenced by our VMI image obtained for S() + Ar collisions.
-This is an excellent starting point for investigating reactive scattering between S() and H2 (including its isotopologues), as well as electronic quenching of S() in collisions with various partners, although it should be noted that the cross sections for these processes are generally expected to be smaller than those for the elastic S() + Ar collisions investigated here.
-Our experimental setup allows for precise control over the collision energy, making it well suited for exploring these processes under well-defined conditions.
+This is an excellent starting point for investigating reactive scattering between S() and H2 (including its isotopologues), as well as electronic quenching of S() in collisions with various partners.
+Since cross sections for such processes could be one to two orders of magnitude smaller than those for elastic scattering, their measurement will be considerably more challenging.
+Nevertheless, the precise control over collision energy provided by our setup would make it well suited to explore these processes under well-defined conditions.
 V Acknowledgements
@@
 101160597).
+VI Data availability
+The data that support the findings of this article are openly available [52].